OpenStudy (anonymous):

solve X+√y=21 y+√x=29

OpenStudy (anonymous):

try and make y = x +something then swap y for that x+ something example y + x = 4 2y + x = 10 so get y alone for one of them. y = 4-x now on the second equation swap y for 4-x 2(4-x) +x = 10 8-2x + x = 10 -2x = 2 x = -1! now that we have x just plug in and solve for y y = 4-x y = 4 - (-1) y = 5!

OpenStudy (anonymous):

\[x + \sqrt{y}=21\] so this tells us that \[x=21-\sqrt{y}\]

OpenStudy (anonymous):

I guess y=25 and x=16 but I want a real solution

OpenStudy (anonymous):

√y=21-x √x=29-y square both term and solve for x and y.

OpenStudy (anonymous):

y=441-42x+x^2 x=841-58x+y^2 what can I do with them?

OpenStudy (anonymous):

if what you did is correct, then everywhere you see y on the second equation plug in what you got for y. y = 441-42x +x^2 right? soooooo x=841-58x+(441-42x+x^2)^2 this is to say y ----^ but you now have a y=something

OpenStudy (anonymous):

and now you can solve for x! once you have x, you can plug that in and solve for y

OpenStudy (anonymous):

i double checked your math and you are on the right track!

OpenStudy (anonymous):

the last EQUATION \[x=841-58(21-x)^{2}+(21-x)^{4}\]